OpenStudy (anonymous):

8th Grade math help asap

OpenStudy (anonymous):

A zoo train ride costs $4 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 27, and the total money collected was $60. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers? 11 children and 16 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60 16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60 11 children and 16 adults Equation 1: a + c = 27 Equation 2: 4a − c = 60 16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a − c = 60

OpenStudy (anonymous):

@kelliegirl33 what about this? lol

OpenStudy (anonymous):

@JK14 you know it?

OpenStudy (anonymous):

I think i can solve it i did this before

OpenStudy (anonymous):

k thank you

OpenStudy (anonymous):

I believe the answer would be the second equation: 16 children and 11 adults Equation 1: a + c = 27 Equation 2: 4a + c = 60

OpenStudy (anonymous):

Question, are you in K12 8th grade?

OpenStudy (anonymous):

flvs 8th grade

OpenStudy (anonymous):

florida virtual school

OpenStudy (anonymous):

ohh cool i was born in florida

OpenStudy (anonymous):

cool !!

OpenStudy (anonymous):

could u help with another question?

OpenStudy (anonymous):

Line A is represented by the equation given below: x + y = 3 What is most likely the equation for line B, so that the set of equations has infinitely many solutions? 3x + 3y = 3 3x + y = 3 x + y = 9 3x + 3y = 9

OpenStudy (anonymous):

sure i'll try

OpenStudy (anonymous):

idk this one at all

OpenStudy (anonymous):

i don't either but i'm guessing it will be 3x+y=3

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

Katie and Pamela put some money into their money boxes every week. The amounts of money (y), in dollars, in their money boxes after certain amounts of time (x), in weeks, are shown by the equations below: Katie y = 30x + 20 Pamela y = 20x + 90 After how many weeks will Katie and Pamela have the same amount of money in their money boxes, and what will that amount be? 6 weeks, $150 6 weeks, $230 7 weeks, $150 7 weeks, $230

OpenStudy (anonymous):

last one i need help with

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@JK14

OpenStudy (anonymous):

I don't know the answer

OpenStudy (anonymous):

30x + 20 = 20x + 90 30x - 20x = 90 - 20 10x = 70 x = 7 y = 30x + 20 y = 20x + 90 y = 30(7) + 20 y = 20(7) + 90 y = 210 + 20 y = 140 + 90 y = 230 y = 230 7 weeks, $230

OpenStudy (anonymous):

I WAS RIGHT :) thats what i put

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

lol Good Job

OpenStudy (anonymous):

A system of equations is shown below: y = 8x − 2 y = 9x − 7 What is the solution to the system of equations? (−5, 38) (−5, −38) (5, 38) (5, −38)

OpenStudy (anonymous):

give me a sec and tell me if im right

OpenStudy (anonymous):

-5, 38?

OpenStudy (anonymous):

close.... 8x - 2 = 9x - 7 8x - 9x = -7 + 2 -x = -5 x = 5 y = 8x - 2 y = 8(5) - 2 y = 40 - 2 y = 38 (5,38)

OpenStudy (anonymous):

dang almost but thanks your a huge help

OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 6m = 4 + 4n Step 1: −6(m) = −6(8 + 2n) [Equation 1 is multiplied by −6.] 6m = 4 + 4n [Equation 2] Step 2: −6m = −48 − 12n [Equation 1 in Step 1 is simplified.] 6m = 4 + 4n [Equation 2] Step 3: −6m + 6m = −48 − 12n + 4n [Equations in Step 2 are added.] Step 4: 0 = −48 − 8n Step 5: n = −6 In which step did the student first make an error? Step 3 Step 5 Step 4 Step 2

OpenStudy (anonymous):

its my 2nd to last one i have

OpenStudy (anonymous):

step 1--correct step 2 -- correct step 3 -- -6m = -48 - 12n 6m = 4 + 4n -------------- 0 = -44 - 8n <====it was added wrong I think that would be step 4....the error occurs in the addition of the 2 equations but possibly step 3 because step 3 is the addition part...step 4 is the result so I am not sure if it is 3 or 4

OpenStudy (anonymous):

k what would u pick?

OpenStudy (anonymous):

probably 4

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

step 4 or D

OpenStudy (anonymous):

step 4 is C isn't it ??

OpenStudy (anonymous):

4 is C yes

OpenStudy (anonymous):

its confusing because the mistake was made during the addition of the 2 equations....so I am a little confused if that would be step 3 or step 4.....you might want to get a second opinion

OpenStudy (anonymous):

idk i trust you

OpenStudy (anonymous):

what about this one? Two equations are given below: m + 5n = 20 m = n − 4 What is the solution to the set of equations in the form (m, n)? (3, 7) (0, 4) (5, 1) (2, 6)

OpenStudy (anonymous):

m = n - 4 so sub in n - 4 in place of m in the other equation m + 5n = 20 n - 4 + 5n = 20 n + 5n = 20 + 4 6n = 24 n = 4 m = n - 4 m = 4 - 4 m = 0 (m,n) = (0,4)

OpenStudy (anonymous):

sorry buddy, but I got to go now....I wish you luck :)

OpenStudy (anonymous):

:( k thank you for your help